# Navigation and Exploration with Active Inference: from Biology to Industry

### v1

@@ Abstract
Keywords:
Autonomous Navigation Active Inference Robotics Topological maps.
@@ 5 Toward realistic settings
Our system is deployed with ROS2 and is modular by design. The AIF-based planner is module, sensor and platform-agnostic and can interact with traditional perception modules and motion planning systems as well as diverse robots (Turtlebot, Turtlebot3 [39] and RosbotXL [16], the used robots and sensors are defined in Appendix Robots). Each component can be independently replaced or refined without disrupting the rest of the pipeline. Future work could incorporate more sophisticated perceptual pipelines (e.g., semantic SLAM or learned embeddings), without requiring changes to the generative model.
@@ 6 Exploration in realistic environments
Overall, our Active Inference framework outperforms or matches the performance of more traditional approaches (Frontiers) and state-of-the-art exploration systems (Gbplanner), with an efficient goal-oriented exploration (goals being the next state to reach). Frontiers falls behind in exploration efficiency as it lacks optimised navigation and requires multiple passes over the same areas. Especially when some small, unreachable areas have unexplored zones without a clear frontier. Gbplanner has been built to be robust in underground passages rather than optimising navigation in large open spaces. Additional experiments validating the dynamic adaptability of our model can be found in appendix Handling Dynamic Obstacles.
@@ Acknowledgements
Figure LABEL:img:ob_move exemplifies this process with an obstacle (e.g. a box) moved between two positions (from position (-1,0) to (-1,-1) over a visited state 3) while the agent explored a mini warehouse [2] presented in Figure 7. As the agent fails to reach state 3 after the movement of the box, it adjusts its internal map to reflect the new reality. The transition probability to that location reduces, and the state ID at this position becomes incorrect (state 1 instead of 3) because the agent cannot correct its belief by moving to that location. A new state (state 20) is created at the former position of the obstacle, and new transitions are established. This change does not affect any of the other existing states.
This display, in a qualitative way, shows how the model effectively adapts to change in the environment. The results displayed in Figure LABEL:img:ob_move is obtained after the agent circled once around the obstacle to update all adjacent states toward the obstructed state.

### v2

@@ 5 Toward realistic settings
Our system is deployed with ROS2 and is modular by design. The AIF-based planner is module, sensor and platform-agnostic and can interact with traditional perception modules and motion planning systems as well as diverse robots (Turtlebot, Turtlebot3 [39] and RosbotXL [16], the used robots and sensors are defined in Appendix 7). Each component can be independently replaced or refined without disrupting the rest of the pipeline. Future work could incorporate more sophisticated perceptual pipelines (e.g., semantic SLAM or learned embeddings), without requiring changes to the generative model.
@@ 6 Exploration in realistic environments
(a) Final map of exploration in a) Amazon simulated warehouse, b) a real-world environment. Coloured points signify visited locations, where the same colour attributions mean the same observation. The thickness of the lines depicts the agent’s believed probability of transitioning between two states given an action.
@@
Overall, our Active Inference framework outperforms or matches the performance of more traditional approaches (Frontiers) and state-of-the-art exploration systems (Gbplanner), with an efficient goal-oriented exploration (goals being the next state to reach). Frontiers falls behind in exploration efficiency as it lacks optimised navigation and requires multiple passes over the same areas. Especially when some small, unreachable areas have unexplored zones without a clear frontier. Gbplanner has been built to be robust in underground passages rather than optimising navigation in large open spaces. Additional experiments validating the dynamic adaptability of our model can be found in appendix 7.
@@ Acknowledgements
This research received funding from the Flemish Government (AI Research Program) under the “Onder-zoeksprogramma Artificiële Intelligentie (AI) Vlaanderen” programme and the Inter-university Microelectronics Centre (IMEC).
@@
(a) Turtlebot3 waffle robot was used in simulation while tests have been conducted with a turtlebot, turtlebot4 and RosbotXL in the real environment
@@
Figure 7(b) exemplifies this process with an obstacle (e.g. a box) moved between two positions (from position (-1,0) to (-1,-1) over a visited state 3) while the agent explored a mini warehouse [2] presented in Figure 7. As the agent fails to reach state 3 after the movement of the box, it adjusts its internal map to reflect the new reality. The transition probability to that location reduces, and the state ID at this position becomes incorrect (state 1 instead of 3) because the agent cannot correct its belief by moving to that location. A new state (state 20) is created at the former position of the obstacle, and new transitions are established. This change does not affect any of the other existing states.
(b) An obstacle was initially placed at position (-1,0) and moved to position (-1,-1) over state 3 during exploration. The failure to reach the state reduces the transition probabilities, represented as thinner black lines between state nodes in the graph. The thickness of the link represents the transition certainty between locations, and each dot colour represents an observation; similar enough observations hold the same colours.
This display, in a qualitative way, shows how the model effectively adapts to change in the environment. The results displayed in Figure 7(b) is obtained after the agent circled once around the obstacle to update all adjacent states toward the obstructed state.